Calculus I Fall 2001 MA113:13-15, 19-21 Russell Brown We have our first exam on Tuesday, 18 September 2001. Below are a few problems to help you review for the exam. In addition to what is listed below, students should be familiar with the correct solutions to homework that was collected and with material presented in lecture.

• §1 #77–82. Students should know careful definition of even and odd functions.

• §2 #5–17.

• Appendix D. # 29, 31, 33, 43, 45, 47, 49. You should know the definitions of all 6 trigonometric functions. The Pythagorean identity, the addition formulas for sine and cosine and symmetry properties (odd/even) of sine and cosine.

• §1.2 #1, 3, 5, 15, 17, 19.

• §1.3 #3, 5, 27, 29, 39.

• §1.4 # 5, 11, 13, 27. Please know the rigorous definition of limit.

• §1.5 # 15, 17, 38, 39, 45, 47. Please know the definition continuous function and the statement of the intermediate value theorem.

• §1.6 #7, 11

• §2.1 #3, 7, 11, 13, 15, 23, 31, 35, 37, 39. Please know the definition of the derivative.